OpenStudy (ilovebmth1234):

To determine whether an element is actually gold, scientists will often determine the density of the element and compare it to the known density of gold (19.3 grams/cubic centimeter, or 19.3 g/cc). Samples less dense might be mixtures of many elements (not pure gold). A sample is found to have a mass of 500 g. The sample is a perfect 2 cm cube. Is this sample gold? Step 1. Identify given and unknown information. Given Information Unknown Information Mass = 500 g Volume = 2 cm x 2 cm x 2 cm = 8 cc Density = ? Step 2. Choose the appropriate formula from the basic scientific equations

OpenStudy (anonymous):

Well, the equation for density is:\[\rho =\frac{ m }{ V }\]where ρ is density; m is mass; and V is volume. It looks like you're given m and V.

OpenStudy (ilovebmth1234):

yeah

OpenStudy (anonymous):

Are you able to solve the problem, now?

OpenStudy (ilovebmth1234):

but how do you divide 500g by 8cc?

OpenStudy (anonymous):

What are the units of density? In this case they are: \[\frac{ g }{ cc }\]In other words, they are grams divided by cc.

OpenStudy (ilovebmth1234):

yes how do you divide two different things like that?

OpenStudy (anonymous):

Simple,you do exact what the equation tell you to do:\[\rho =\frac{ m }{ V }=\frac{ 500g }{ 8cc }=62.5 g/cc\]That most certainly not gold--or any known element for that matter--because the density is far too high. Are you sure the values in the problem are correct?

OpenStudy (ilovebmth1234):

yep